# Error mitigation for logical circuits using decoder confidence

### v1

@@ Abstract
Fault-tolerant quantum computers use decoders to monitor for errors and find a plausible correction. A decoder may provide a decoder confidence score (DCS) to gauge its success. We adopt a swim distance DCS, computed from the shortest path between syndrome clusters. By contracting tensor networks, we compare its performance to the well-known complementary gap and find that both reliably estimate the logical error probability (LEP) in a decoding window. We explore ways to use this to mitigate the LEP in entire circuits. For shallow circuits, we just abort if any decoding window produces an exceptionally low DCS: for a distance-13 surface code, rejecting a mere $0.1\%$ of possible DCS values improves the entire circuit’s LEP by more than 5 orders of magnitude. For larger algorithms comprising up to trillions of windows, DCS-based rejection remains effective for enhancing observable estimation. Moreover, one can use DCS to assign each circuit’s output a unique LEP, and use it as a basis for maximum likelihood inference. This can reduce the effects of noise by an order of magnitude at no quantum cost; methods can be combined for further improvements.
@@ II.1 The Role of a Decoder
More formally, the decoding problem involves the decoding graph $G=(V_{\textnormal{d}}\cup V_{\textnormal{b}},E)$, with the set $V_{\textnormal{d}}$ of detectors and the set $V_{\textnormal{b}}$ of boundary nodes such that $V_{\textnormal{d}}\cap V_{\textnormal{b}}=\varnothing$. Each boundary node in $V_{\textnormal{b}}$ represents a unique boundary of the surface code (so for a memory experiment, $|V_{\textnormal{b}}|=2$). An example of how $G$ relates to its code is shown in Fig. 1.
@@
The logical error probability (LEP), denoted $p_{\textnormal{L}}$, is the probability that the decoder is not successful. The log success odds (or log-likelihood ratio of success) is given by
@@ II.2 Decoder Confidence Scores
Given $G$ and $\mathbb{S}$, the DCS is an estimator of some monotonic function $\alpha(\lambda)$ of $\lambda$.
@@ III Decoder confidence scores are an excellent indicator of logical error risk
Figure 5: Single-decoding-window statistics. (a) Histograms of DCS (swim distance) values, normalized by the number of shots. Under circuit-level noise, $10^{8}$ memory experiments were run for all code distances. The measured DCS values are grouped in 50 bins of equal linear width. (b) The log success odds conditioned on the DCS value, plotted against the DCS value. The upper and lower bounds on the Bernoulli variable $p_{L}$ are given by the Wilson score interval. Only datapoints with $p_{L}>0$ are shown.
@@ IV Decoder confidence score distributions are rich in information
In this section, we explore statistical properties of the DCS distribution which make it a valuable resource for error mitigation at both the single-window level, in Section IV.1, and the full-application level, in Section IV.2. Throughout the rest of this paper, we focus on the swim distance DCS, and we will henceforth generally refer to it as ‘the DCS’. We will refer to single-window LEPs and logical errors by lowercase symbols ($p_{\textnormal{L}},x$) and to their multi-window counterparts by uppercase symbols ($P_{\textnormal{L}},X$). We emulate circuits using Stim [28] and assume circuit-level noise with a physical error rate $p_{\text{phys}}=10^{-3}$. Our approach to estimating DCS performance in the context of full logical circuits is based on allocating a suitable number of decoding windows. We do not adjust according to e.g. the specific decoding needs of logical gates, leaving that for future studies.
@@ IV.1 Single-decoding-window statistics
We emulated memory experiments in the rotated surface code for odd code distances $d\in\{5,\dots,19\}$; these assess how well one logical qubit stays free from logical error after $d$ stabiliser measurement cycles. We decode each experiment in a single window and calculate the DCS for that window. The distribution of these values is shown in Fig. 5a. Each bin in this histogram contains a number of failed experiments (where a logical error occurred) and successful (i.e. error-free) experiments, which we use to calculate an empirical log success odds from Eq. 1. Figure 5b plots this empirical log success odds against the DCS. By fitting a straight line to this data, we obtain a calibration curve that assigns a LEP $p_{\textnormal{L}}$ to each DCS value. These curves can be used in future experiments for the same code distance and noise model.
@@ IV.2 Many-decoding-window statistics
Figure 6: Equivalence of different code distances. (a) Histogram of the calibrated LEPs, normalized by the number of shots, for various numbers of decoding windows in the circuit. The LEPs are grouped in 50 bins of equal logarithmic width. At fixed code distance $d$, for circuits comprising $N$ windows, we define $f_{d,N}(p_{\textnormal{L}})$ as the average number of constituent windows whose LEP lies in the bin centred on $p_{\textnormal{L}}$. The case $N=1$ (viridis colourmap) corresponds to the probability mass function of LEPs for the single-window memory experiments. When more windows are added (magma colourmap), all bins gain population at the same rate: $f_{d,N}(p_{\textnormal{L}})=Nf_{d,1}(p_{\textnormal{L}})$. It is possible to adjust NN such that circuits defined by different $(N,d)$ pairs have the same mean $P_{\textnormal{L}}$; in this case, the seemingly linear segments of the scaled histograms overlap, as indicated by the upward arrows. (b) The calculated LEP (solid) and observed LER (dashed) in a cohort of single-window decoding instances when a fraction of highest-$p_{\textnormal{L}}$ shots are discarded. The curves for $d\in\{{11,13}\}$ were obtained with $10^{9}$ shots.
To simplify analysis, we assume $(p_{\textnormal{L}},x)$ for each window is independent; this simplification is discussed in Section VII. In this case, the total LEP $P_{\textnormal{L}}$ for an experiment comprising $N$ decoding windows, each with LEP $p_{{\textnormal{L}}_{i}}$, is the probability an odd number of them were wrongly decoded:
$P_{\textnormal{L}}=\frac{1}{2}\left[1-\prod_{i=1}^{N}(1-2p_{\textnormal{L},i})\right].$ (3)
@@
To investigate this, we generalise Fig. 5a to circuits of $N\geq 1$ decoding windows and plot the result in Fig. 6a, where the $x$-axis has additionally been converted to a LEP via the calibration curve. Each single-window LEP histogram is scaled by a factor $N$ such that the mean LEP of an $N$-window circuit equals $1\text{\times}{10}^{-2}$. The evident overlap of the scaled histograms for $p_{\textnormal{L}}>10^{-7}$ means that the same number of high-$p_{\textnormal{L}}$ windows will be detected for different $(N,d)$ pairs. These rare, high-risk events are precisely the ones which make the dominant contribution to the overall $N$-window circuit LEP $P_{\textnormal{L}}$. Because the non-overlapping parts of the scaled histograms do not contribute much to the overall LEP, the $P_{\textnormal{L}}$ distribution for a numerically intractably large pair $(N,d)$ can be accurately inferred from the $P_{\textnormal{L}}$ distribution for a smaller, testable pair $(N^{\prime},d^{\prime})$ yielding the same mean LEP. The fact that the low-$p_{\textnormal{L}}$ windows bring little contribution to $P_{\textnormal{L}}$ can be seen from Fig. 6b, where we consider the scenario of discarding a fraction of the single-window shots with the highest $p_{\textnormal{L}}$ values. The figure shows both the calculated LEP and the empirically observed LER; these lines agree well. For $d=13$ it is possible to improve the LER by one (five) orders of magnitude by discarding a fraction $1\text{\times}{10}^{-7}$ ($1\text{\times}{10}^{-3}$) of the total shots.
@@ V.2 DCS-Based Maximum Likelihood Estimation
We consider estimating the expectation value of an observable from $M$ noisy runs of the same quantum circuit comprising $N$ decoding windows. Each run $j\in\{{1,\dots,M}\}$ has an observed noisy outcome $z_{j}\in\{{-1,+1}\}$ and a total LEP $P_{\textnormal{L},j}$ calculated from the DCS. Section C.2 details how we emulate these runs. A straightforward estimator is the unweighted average: $\langle Z_{\textrm{unmitigated}}\rangle=\frac{1}{M}\sum_{j=1}^{M}z_{j}.$ However, this can be improved using the data $\{{P_{\textnormal{L},j}}\}_{j}$ as follows.
@@
where $\operatorname{pr}(\text{logical statement})$ is the probability the logical statement is true. Let the underlying noiseless outcomes from the logical circuit be $\{\tilde{z}_{j}\}_{j=1}^{M}$, all having the same mean, $\langle Z\rangle_{\text{th}}$, such that $\theta\equiv\operatorname{pr}(\tilde{Z}_{j}=+1)=\frac{1}{2}(1+\langle Z\rangle_{\text{th}})$. We can express the probability to observe a particular outcome in terms of conditional probabilities on the noiseless outcomes:
$\operatorname{pr}(Z_{j}=z_{j})=\sum_{\tilde{z}_{j}}\operatorname{pr}(Z_{j}=z_{j}|\tilde{Z}_{j}=\tilde{z}_{j})\operatorname{pr}(\tilde{Z}_{j}=\tilde{z}_{j})$ (5)
Generally, noise corrupts the true outcomes in a non-trivial way, defined by a corruption function $q$ which depends on the logical error probability: $q(P_{\textnormal{L},j},\tilde{z}_{j})\equiv\operatorname{pr}(Z_{j}=+1|\tilde{Z}_{j}=\tilde{z}_{j})$. Therefore:
@@
To tackle imperfect LEP–DCS calibration, we use a two-parameter fit: we replace $P_{\textnormal{L},j}$ by $\eta P_{\textnormal{L},j}$ in Eq. 6 and also optimise the parameter $\eta$. This accounts for an overall rescaling of the original LEP distribution, to ensure the mean LEP matches the true LER.
@@ V.3 Numerics
Figure 7: Estimating $\langle Z\rangle_{\text{th}}=0.8$ on a noisy quantum computer, starting from a $7.7(2)\%$ LER in absence of error mitigation. (a,c) Output distributions of four expectation value estimators, obtained from $N_{\text{rep}}=2\times 10^{5}$ repetitions. Dashed curves are normal distributions with the same mean and variance as their respective datasets, which are shown as dots. The noiseless estimator corresponds to a noiseless quantum computer whose outcome variation is caused solely by statistical Bernoulli fluctuations. The circuit discard fraction is $50\%$. (b,d) The performance of the four estimators plotted against resource overhead. MSPE $\sum_{i=1}^{N_{\text{rep}}}|\langle Z\rangle_{i}-\langle Z\rangle_{\text{th}}|^{2}/N_{\text{rep}}$ is the mean squared prediction error. The overhead calculation is explained in Appendix D.
We next evaluate the methods from the previous two subsections on emulated results of an expectation value estimation task using a distance-11 surface code. Three main observations emerge: 1) both methods significantly lower the total error; 2) MLE gives the highest relative improvement over pure abort when applied to unfiltered data; 3) both methods give higher relative improvements as the number of shots increases. We also construct an approximate model for the imperfect mapping between the DCS and the success log odds for a distance-19 surface code and assess the performance of the abort protocol under this model.
We set to estimate $\langle Z\rangle_{\text{th}}=0.8$ from $M_{0}\in\{$1\text{\times}{10}^{2}$,$1\text{\times}{10}^{4}$\$\} accepted shots with a $d=11$ surface code, in a circuit comprising $N\approx 2.38\times 10^{5}$ decoding windows. We consider a range of abort rates, each of which gives rise to a different LEP distribution for the $N$-window circuit. Appendix C describes how we generate the datasets from empirical data. Appendix D explains how to calculate the time overhead given the abort rate.
In Fig. 7a,c, we show the distribution of outputs from four different estimators. In Fig. 7b,d, we choose the mean squared prediction error (MSPE) as a performance metric as it penalizes both bias and variance in the estimated expectation values. Other metrics could be selected based on the task at hand, such as the absolute mean bias (shown in Fig. G.1) or the probability to obtain an estimate within a desired range.
@@
Figure 8: Contrasting the at-scale performance of the DCS (blue) versus the ideal but impractical-to-compute log odds (gray). In all cases the chosen DCS is the swim distance. In panel (a) the green dashed line shows the distribution of measured DCS values for decoding windows of a distance-19 surface code. An assumed probability curve for the log odds is shown in gray, and in blue an implied DCS distribution derived by redistributing the log odds values via a Gaussian with standard deviation $\Delta=1$. We see a good match between green and blue. Using the distributions in (a) we generate sets of $2.8\text{\times}{10}^{9}$ scores, representing all decoding windows in a large circuit; an example distribution is shown in (b). From such a set we infer the probability $P_{\textnormal{L}}$ of error in the circuit; repeating this 105 times we obtain distribution (c) with mean $P_{\textnormal{L}}=0.153$. We note that relying on the DCS means we cannot identify very high or very low risk circuits, as we would with access to the log odds. Panel (d) shows the factor by which we reduce error rates using abort criteria, at the cost of increased quantum processor time. Panel (e) is the equivalent of Fig. 7b,d with the same meaning for lines; here we use 200 shots (between the 100 & 1000 Fig. 7 scenarios).

### v2

@@ Abstract
Fault-tolerant quantum computers use decoders to monitor for errors and find a plausible correction. A decoder may provide a decoder confidence score (DCS) to gauge its success. We adopt a swim distance DCS, computed from the shortest path between syndrome clusters. By contracting tensor networks, we compare its performance to the well-known complementary gap and find that both reliably estimate the logical error probability (LEP) in a decoding window. We explore ways to use this to mitigate the LEP in entire circuits. For shallow circuits, we just abort if any decoding window produces an exceptionally low DCS: for a distance-13 surface code, rejecting a mere $0.1\%$ of possible DCS values improves the entire circuit’s LEP by more than 5 orders of magnitude. For larger algorithms comprising up to trillions of windows, DCS-based rejection remains effective for enhancing observable estimation. Moreover, one can use the DCS to assign each circuit’s output a unique LEP, and use it as a basis for maximum likelihood estimation. This can reduce the effects of noise by an order of magnitude at no quantum cost; methods can be combined for further improvements.
@@ II.1 The Role of a Decoder
[…]
@@
The logical error probability (LEP), denoted $p_{\textnormal{L}}$, is the probability that the decoder is not successful, conditioned on $\mathbb{S}$. The log success odds (or log-likelihood ratio of success) is given by
@@ II.2 Decoder Confidence Scores
Given $G$ and $\mathbb{S}$, the DCS is an estimator of some monotonic function $\alpha(\lambda)$ of $\lambda$. We denote its value as $\phi$.
@@ III Decoder Confidence Scores are an Excellent Indicator of Logical Error Risk
Figure 5: Single-decoding-window statistics. (a) Histograms of the DCS (swim distance) value $\phi$, normalised by number of shots. Under circuit-level noise, $10^{8}$ memory experiments were run for all code distances. The $\phi$ values are grouped in 50 bins of equal linear width. (b) The log success odds conditioned on $\phi$, plotted against $\phi$. The upper and lower bounds on the Bernoulli variable $p_{\textnormal{L}}$ are given by the Wilson score interval. Only datapoints with $p_{\textnormal{L}}>0$ are shown.
@@ IV Decoder Confidence Score Distributions are Rich in Information
Figure 6: Equivalence of different code distances. (a) Histogram of the calibrated LEPs, normalised by the number of shots, for various numbers of decoding windows in the circuit. The LEPs are grouped in 50 bins of equal logarithmic width. At fixed code distance $d$, for circuits comprising $N$ windows, we define $f_{d,N}(p_{\textnormal{L}})$ as the average number of constituent windows whose LEP lies in the bin centred on $p_{\textnormal{L}}$. The case $N=1$ (viridis colourmap) corresponds to the normalized LEP histogram for single-window memory experiments. When more windows are added (magma colourmap), all bins gain population at the same rate: $f_{d,N}(p_{\textnormal{L}})=Nf_{d,1}(p_{\textnormal{L}})$. It is possible to adjust NN such that circuits defined by different $(N,d)$ pairs have the same mean $P_{\textnormal{L}}$; in this case, the seemingly linear segments of the scaled histograms overlap, as indicated by the upward arrows. (b) The calculated mean LEP (solid) and observed LER (dashed) in a cohort of single-window decoding instances when a fraction of highest-$p_{\textnormal{L}}$ shots are discarded. The curves for $d\in\{{11,13}\}$ were obtained with $10^{9}$ shots.
In this section, we explore statistical properties of the DCS distribution which make it a valuable resource for error mitigation at both the single-window level, in Section IV.1, and the full-application level, in Section IV.2. Throughout the rest of this paper, we focus on the swim distance DCS, and we will henceforth generally refer to it as ‘the DCS’ and its value as $\phi$. We will refer to single-window LEPs and logical errors by lowercase symbols ($p_{\textnormal{L}},x$) and to their multi-window counterparts by uppercase symbols ($P_{\textnormal{L}},X$). We emulate circuits using Stim [28] and assume circuit-level noise with a physical error rate $p_{\text{phys}}=10^{-3}$. Our approach to estimating DCS performance in the context of full logical circuits is based on allocating a suitable number of decoding windows. We do not adjust according to e.g. the specific decoding needs of logical gates, leaving that for future studies.
@@ IV.1 Single-Decoding-Window Statistics
We emulated memory experiments in the rotated surface code for odd code distances $d\in\{5,\dots,19\}$; these assess how well one logical qubit stays free from logical error after $d$ stabiliser measurement cycles. We decode each experiment in a single window and calculate $\phi$ for that window. The distribution of $\phi$ is shown in Fig. 5a. Each bin in this histogram contains a number of failed experiments (where a logical error occurred) and successful (i.e. error-free) experiments, which we use to calculate an empirical log success odds from Eq. 1. Figure 5b plots this empirical log success odds against $\phi$. By fitting a straight line to this data, we obtain a calibration curve that assigns a LEP $p_{\textnormal{L}}$ to each $\phi$ value. These curves can be used in future experiments for the same code distance and noise model.
@@ IV.2 Many-Decoding-Window Statistics
To simplify analysis, we assume $(p_{\textnormal{L}},x)$ for each window is independent; this simplification is discussed in Section VII. In this case, the total LEP $P_{\textnormal{L}}$ for an experiment comprising $N$ decoding windows, each with LEP $p_{\textnormal{L},i}$, is the probability an odd number of them were wrongly decoded:
$P_{\textnormal{L}}=\frac{1}{2}\bigg[{1-\prod_{i=1}^{N}(1-2p_{\textnormal{L},i})}\bigg].$ (3)
@@
To investigate this, we generalise Fig. 5a to circuits of $N\geq 1$ decoding windows and plot the result in Fig. 6a, where the $x$-axis has additionally been converted to a LEP via the calibration curve. Each single-window LEP histogram is scaled by a factor $N$ such that the mean LEP of an $N$-window circuit equals $1\text{\times}{10}^{-2}$. The evident overlap of the scaled histograms for $p_{\textnormal{L}}>10^{-7}$ means that the same number of high-$p_{\textnormal{L}}$ windows will be detected for different $(N,d)$ pairs. These rare, high-risk events are precisely the ones which make the dominant contribution to the overall $N$-window circuit LEP $P_{\textnormal{L}}$.
Remark 1.
Because the non-overlapping parts of the scaled histograms do not contribute much to the overall LEP, the $P_{\textnormal{L}}$ distribution for a numerically intractably large pair $(N,d)$ can be accurately inferred from the $P_{\textnormal{L}}$ distribution for a smaller, testable pair $(N^{\prime},d^{\prime})$ yielding the same mean LEP.
The fact that the low-$p_{\textnormal{L}}$ windows bring little contribution to $P_{\textnormal{L}}$ can be seen from Fig. 6b, where we consider the scenario of discarding a fraction of the single-window shots with the highest $p_{\textnormal{L}}$ values. The figure shows both the calculated LEP and the empirically observed LER; these lines agree well. For $d=13$ it is possible to improve the LER by one (five) orders of magnitude by discarding a fraction $1\text{\times}{10}^{-7}$ ($1\text{\times}{10}^{-3}$) of the total shots.
@@ V.2 DCS-Based Maximum Likelihood Estimation
We consider estimating the expectation value of an observable from $M$ noisy runs of the same quantum circuit comprising $N$ decoding windows. Each run $j\in\{{1,\dots,M}\}$ has an observed noisy outcome $z_{j}\in\{{-1,+1}\}$ and a total LEP $P_{\textnormal{L},j}$ calculated from $\phi$. Section C.2 details how we emulate these runs. A straightforward estimator is the unweighted average: $\langle Z_{\textrm{unmitigated}}\rangle=\frac{1}{M}\sum_{j=1}^{M}z_{j}.$ However, this can be improved using the data $\{{P_{\textnormal{L},j}}\}_{j}$ as follows.
@@
where $\operatorname{pr}(\text{logical statement})$ is the probability the logical statement is true. Let the underlying noiseless outcomes from the logical circuit be $\{\tilde{z}_{j}\}_{j=1}^{M}$, all having the same mean, $\langle Z\rangle_{\text{th}}$, such that $\theta:=\operatorname{pr}(\tilde{Z}_{j}=+1)=\frac{1}{2}(1+\langle Z\rangle_{\text{th}})$. We can express the probability to observe a particular outcome in terms of conditional probabilities on the noiseless outcomes:
$\operatorname{pr}(Z_{j}=z_{j})=\sum_{\tilde{z}_{j}}\operatorname{pr}(Z_{j}=z_{j}|\tilde{Z}_{j}=\tilde{z}_{j})\operatorname{pr}(\tilde{Z}_{j}=\tilde{z}_{j}).$ (5)
Generally, noise corrupts the true outcomes in a non-trivial way, defined by a corruption function $q$ which depends on the logical error probability: $q(P_{\textnormal{L},j},\tilde{z}_{j}):=\operatorname{pr}(Z_{j}=+1|\tilde{Z}_{j}=\tilde{z}_{j})$. Therefore:
@@
To tackle imperfect $P_{\textnormal{L}}$$\phi$ calibration, we use a two-parameter fit: we replace $P_{\textnormal{L},j}$ by $\eta P_{\textnormal{L},j}$ in Eq. 6 and also optimise the parameter $\eta$. This accounts for an overall rescaling of the original LEP distribution, to ensure the mean LEP matches the true LER.
@@ V.3 Numerics
Figure 7: Estimating $\langle Z\rangle_{\text{th}}=0.8$ using a distance-11 surface code, starting from a $7.7(2)\%$ LER in absence of error mitigation. (a,c) Output distributions of four expectation value estimators, obtained from $N_{\text{rep}}=2\times 10^{5}$ repetitions. Dashed curves are normal distributions with the same mean and variance as their respective datasets, which are shown as dots. The noiseless estimator corresponds to a noiseless quantum computer whose outcome variation is caused solely by statistical Bernoulli fluctuations. The circuit discard fraction is $50\%$. (b,d) The performance of the four estimators plotted against resource overhead. MSPE is the mean squared prediction error, defined in Eq. 7. The multiplicative time cost calculation is explained in Appendix D.
We next evaluate the methods from the previous two subsections on emulated results of an expectation value estimation task using a distance-11 surface code (Section V.3.1). Three main observations emerge: 1) both methods significantly lower the total error; 2) MLE gives the highest relative improvement over pure abort when applied to unfiltered data; 3) both methods give higher relative improvements as the number of shots increases. We also construct an approximate model for the imperfect mapping between $\phi$ and $\lambda$ for a distance-19 surface code and assess the performance of the abort protocol under this model (Section V.3.2).
V.3.1 Distance-11 Surface Code
We set to estimate $\langle Z\rangle_{\text{th}}=0.8$ from $M_{0}\in\{$1\text{\times}{10}^{2}$,$1\text{\times}{10}^{4}$\$\} accepted shots with a $d=11$ surface code, in a circuit comprising $N\approx 2.38\times 10^{5}$ decoding windows. We consider a range of abort rates, each of which gives rise to a different LEP distribution for the $N$-window circuit. Appendix C describes how we generate the datasets, and Appendix D explains how we calculate the multiplicative time cost given the abort rate.
In Fig. 7a,c, we show the distribution of outputs from four different estimators. Figure 7b,d shows how they perform as the resource overhead varies. We choose the mean squared prediction error (MSPE):
$\textnormal{MSPE}:=\frac{1}{N_{\text{rep}}}\sum_{i=1}^{N_{\text{rep}}}|\langle Z\rangle_{i}-\langle Z\rangle_{\text{th}}|^{2}$ (7)
as a performance metric as it penalises both bias and variance in the estimated expectation values. Other metrics could be selected based on the task at hand, such as the absolute mean bias (explored in Appendix G) or the probability to obtain an estimate within a desired range.
@@ V.3.1 Distance-11 Surface Code
Figure 8: Contrasting the at-scale performance of the DCS (blue) versus the ideal but impractical-to-compute log odds $\lambda$ (gray). In all cases the chosen DCS is the swim distance. In panel (a) the green dashed line shows the fitted probability density function of measured DCS values for decoding windows of a distance-19 surface code, by binning the calibration data into 500 bins. An assumed probability curve for $\lambda$ is shown in gray, and in blue an implied DCS distribution derived by redistributing $\lambda$ via a Gaussian with standard deviation $\Delta=1$. We see a good match between green and blue. Using the distributions in (a) we generate sets of $2.8\text{\times}{10}^{9}$ scores, representing all decoding windows in a large circuit; an example distribution is shown in (b). From such a set we infer the probability $P_{\textnormal{L}}$ of error in the circuit; repeating this 105 times we obtain distribution (c) with mean $P_{\textnormal{L}}=0.153$. We note that relying on the DCS means we cannot identify very high or very low risk circuits, as we would with access to $\lambda$. Panel (d) shows the factor by which we reduce error rates using abort criteria, at the cost of increased quantum processor time. Panel (e) is the equivalent of Fig. 7b,d with the same meaning for lines, but with an intermediate number of shots: 200.
V.3.2 Distance-19 Surface Code